find two numbers whose sum is 10 and whose product is 32. use the quadratic formula to solve
Answer by stanbon(57222)

Find two numbers whose sum is 10 and whose product is 32. use the quadratic formula to solve
--------------------------------------------------------------------------------
Equations:
x+y = 10
xy = 32
---
Substitute for y and solve for "x":
x(10-x) = 32
---
-x^2+10x-32 = 0
---
x^2-10x+32 = 0
-----
x = [10 +- sqrt[100-4*32]]/2
----
x = [10 +- sqrt(-28)]/2
----
x = [10 +- 2isqrt(7)]/2
---
x = (5+isqrt(7)) or x = (5-isqrt(7))
-----
Cheers,
Stan H.